OpenStudy (anonymous):

Which value for x makes the open sentence true? 5 + 4 • x = 32 + x2 A. 3 B. 4 C. 5 D. 2

OpenStudy (anonymous):

find the X in the above equation and that should be your solution

OpenStudy (anonymous):

5+4*3=32+3^2 5+12=32+9 17=41 5+4*4=32+4^2 5+16=32+16 21=48 5+4*5=32+5^2 5+20=32+25 25=57 5+4*2=32+2^2 5+8=32+4 13=36 non of those are actually equal

OpenStudy (anonymous):

5+4x=32+x^2 minus 32 4x-27=x^2 divide by x 4-27=x x=-23 x would have to be -23 for that equation to work

OpenStudy (anonymous):

I appreciate the hard work and the effort you put forth. But the answer is 3, 4, 5, or 2 :c